# The Coherence Trap: When MLLM-Crafted Narratives Exploit Manipulated Visual Contexts

### v1

@@ Abstract
The detection and grounding of multimedia manipulation has emerged as a critical challenge in combating AI-generated disinformation. While existing methods have made progress in recent years, we identify two fundamental limitations in current approaches: (1) Underestimation of MLLM-driven deception risk: prevailing techniques primarily address rule-based text manipulations, yet fail to account for sophisticated misinformation synthesized by multimodal large language models (MLLMs) that can dynamically generate semantically coherent, contextually plausible yet deceptive narratives conditioned on manipulated images; (2) Unrealistic misalignment artifacts: currently focused scenarios rely on artificially misaligned content that lacks semantic coherence, rendering them easily detectable. To address these gaps holistically, we propose a new adversarial pipeline that leverages MLLMs to generate high-risk disinformation. Our approach begins with constructing the MLLM-Driven Synthetic Multimodal (MDSM) dataset, where images are first altered using state-of-the-art editing techniques and then paired with MLLM-generated deceptive texts that maintain semantic consistency with the visual manipulations. Building upon this foundation, we present the Artifact-aware Manipulation Diagnosis via MLLM (AMD) framework featuring two key innovations: Artifact Pre-perception Encoding strategy and Manipulation-Oriented Reasoning, to tame MLLMs for the MDSM problem. Comprehensive experiments validate our framework’s superior generalization capabilities as a unified architecture for detecting MLLM-powered multimodal deceptions.
@@ 1 Introduction
[…]
figureComparison between $\text{DGM}^{4}$ and MDSM (ours). (a) $\text{DGM}^{4}$ separates manipulation of two modalities, using rule-based text editing lacking image integration, making samples easily distinguishable. (b) We use fine-grained facial editing to guide MLLMs in generating contextually consistent, deceptive texts, making samples harder to detect.
@@
Generative techniques have become a transformative force in artificial intelligence[44, 48, 31, 10, 4], demonstrating remarkable adaptability across various domains and modalities. These advancements, while enriching multimedia content, also pose significant challenges to information security. In the media industry in particular, maliciously fake content manipulated by such models can profoundly mislead audiences [51, 37]. The unchecked spread of fake media has already negatively affected political, financial, and other sectors[6, 32, 35], gradually becoming a major social issue[27].
@@
To confront the above issues, many attempts have been made to detect fake multimedia and ground fine-grained forgery components[21, 38, 52, 49]. While various fake news scenarios have been explored, including Luo et al.’s[24] investigation of out-of-context social image-text pairs and Shao et al.’s[38] work on detecting randomly tampered regions or words, our analysis reveals two critical limitations in existing research: 1) Neglect of emerging risks from MLLMs: Current paradigms predominantly address rule-based text manipulation, overlooking the sophisticated linguistic capabilities of modern multi-modal large language models (MLLMs). MLLM-generated text exhibits superior fluency and contextual coherence, significantly increasing deception potential and public susceptibility. 2) Semantic misaligned artifacts. Most methodologies independently manipulate visual and textual elements, producing semantically discordant multimedia outputs. This misalignment not only renders manufactured disinformation too simplistic to effectively deceive the public, but also fails to replicate real-world adversarial behavior, as sophisticated attackers typically maintain meticulous visual-textual consistency to maximize manipulative impact. These limitations render the multi-modal disinformation scenarios considered in existing works insufficiently realistic.
@@
To address these weaknesses, we take MLLM into consideration and focus on detecting the semantic-aligned manipulation. We first construct the MLLM-Driven Synthetic Multi-modal (MDSM) dataset, which manipulates image and text in a coordinated fashion using MLLM. For the image manipulation, we consider the typical Face Swap and Face Attribute editing. For text, We innovatively guide MLLM to generate modality-aligned yet misleading fake narratives using image editing metadata. As shown in Fig. 1(b), after replacing Donald Trump’s face with Micheál Martin’s, we use the swapped name, Micheál Martin, to guide MLLM in generating text, ensuring that the named entity in the text aligns with the image. Following this strategy, we construct over 441k sample pairs.
@@
The alignment of modalities and the authentic texts from MLLMs pose significant challenges for the detection of manipulated media. First, the strategy of perceiving inconsistencies between images and text through contrastive learning, as employed by prior works[52, 38], is ineffective in MDSM where images and text are well-matched already. Merely observing aligned image-text pairs is inadequate for reliable detection. Consequently, external clues and contextual knowledge are essential. Second, existing architectures like ASAP[52] and HAMMER[38], which feature multiple detection and grounding heads, are complex and lack generalizability to unseen media sources. To address these challenges, we propose Artifact-aware Manipulation Diagnosis via MLLM (AMD), which leverages MLLMs’ comprehensive understanding of real-world multimedia and their ability to provide unified textual outputs. And AMD generates detection and grounding results in a coherent, text-based format, offering a more intuitive and generalized solution.
In summary, we highlight our contributions of this paper as follows:
@@
  • Comprehensive evaluations validate AMD’s effectiveness and generalization capability, outperforming existing methods while maintaining parameter efficiency.
@@ 2 Related Work
Deepfake Detection. With the rapid advancement of generative models and the explosive growth of generated data, Deepfake detection technology has also progressed swiftly, enabling it to address the information risks posed by counterfeit data. Historically, Deepfake detection has been categorized into unimodal and multimodal detection approaches. Unimodal detection includes both image-based [53, 19] and text-based [40, 12] methods, all of which have achieved strong intra-domain performance. Recently, with the rise of Multimodal Large Language Models (MLLMs), multimodal Deepfake detection has garnered increasing attention [21, 38, 22]. Among them, HAMMER [38] combines contrastive learning to build a robust detection model that not only classifies manipulation types but also grounds manipulation locations. However, HAMMER does not address the enhancement of cross-domain performance. FKA-Owl [22], which is based on MLLM, incorporates more world knowledge to improve the model’s cross-domain performance. Despite this, FKA-Owl lacks grounding functionality and, due to the incorporation of MLLM, has become particularly cumbersome.
@@
Multi-Modal Large Language Model. In recent years, Multi-Modal Large Language Models (MLLMs) have emerged as a crucial technology for understanding and reasoning across multiple modalities, particularly text and images. By extending the capabilities of Large Language Models (LLMs) to incorporate visual inputs, these models have demonstrated outstanding performance in tasks such as image captioning and visual question answering. CLIP [33] and ALIGN [15] leveraged contrastive learning to align visual and textual representations, enabling efficient zero-shot vision-language understanding. Subsequently, models such as Flamingo [1] and BLIP-2 [18] have introduced vision-language transformers, integrating pre-trained LLMs with vision encoders to enhance cross-modal reasoning and generative capabilities. More recently, GPT-4V [28] and Florence-2 [4] have significantly enhanced the potential of MLLMs in tackling complex multi-modal tasks by leveraging a more efficient framework and larger-scale pre-training data. A key advantage of MLLMs is their ability to acquire extensive world knowledge during large-scale pretraining, which significantly enhances their reasoning capabilities in downstream tasks.
In this work, we fully leverage the intrinsic world knowledge of MLLMs to enhance the robustness of multi-modal manipulation detection and improve inference capabilities in unknown scenarios.
@@ 3 MDSM Dataset Construction
Figure 1: Construction pipeline of MDSM. (a) Face Swap and Text Fabrication steps, where Trump’s face is swapped with Musk’s. (b) Face Attribute and Text Fabrication steps, where Hammond’s expression is edited to anger. (c) Leveraging MLLM to analyze manipulated facial identities in the image and craft tailored prompts to guide it in generating forgery texts.
Fig. 1 illustrates the MDSM construction pipeline. Multi-modal news samples are first collected from diverse sources, alongside a Celebrity Head-shot reference dataset. These resources then drive a two-stage synthesis process: 1) advanced image editing models generate visual manipulations, 2) MLLMs produce semantically aligned textual narratives. We elaborate on these components below.
@@ 3.1 Multi-Modal Media Source Collection
We use the GoodNews [5], VisualNews [11], and N24News [54] datasets as the original datasets $\mathcal{D}_{O}$, which consist of image-text pairs sourced from various real-world news outlets. Given the significant influence of human-centric news among various forms of multi-modal media, we focus on human-centric data for MDSM. $\mathcal{D}_{O}$ is filed by detecting faces in images with Dlib [17] and identifying person names in texts with BERT [8]. Only pairs containing both faces and named entities are kept in the source pool $\mathcal{D}_{S}=\{p_{s}|p_{s}=(I_{s},T_{s})\}$ for manipulation.
3.2 Celebrity Head-shot Dataset
Before detailing the construction of MDSM, we first collect the Celebrity Head-shot Dataset $\mathcal{D}_{C}$, which contains over 29k pairs of head-shot images and corresponding names to facilitate the aligned manipulation of multi-modal media.
@@ 3.3 Multi-Modal Media Manipulation
In the image modality, two main attacks, Face Swap (FS) and Face Attribute (FA), are employed. For the text modality, we utilize advanced MLLMs to generate semantic-aligned texts for the images.
$\triangleright$ Face Swap Manipulation. Face swap is a critical tool for attackers to forge images of public figures and politicians, posing threats to societal security (Fig. 1a). We use two representative face swap methods, SimSwap [7] and e4s [23], to perform such manipulations. We prioritize modifying larger faces to target the primary subject in the image (Fig. 2c). Given a source image $I_{s}$, we randomly choose one of the two methods to replace the largest face $F_{s}$ in $I_{s}$ with a face $F_{c}$ from $\mathcal{D}_{C}$, generating a manipulated face swap sample $I_{fs}$. The bounding box $y_{box}=\{x_{1},y_{1},x_{2},y_{2}\}$ of the swapped face is stored as an annotation, and the name $N_{c}$ of $F_{c}$ is recorded for text fabrication.
Table 1: Comparison of the proposed MDSM with existing misinformation datasets, where MM Det., Text Det., Man. Type Det., and Im. GD stand for Multi-media Detection, Text Detection, Manipulation Type Detection and Image Grounding.
Datasets Samples Modality Tasks Semantic MLLM
Text Image MM Det. Text Det. Man. Type Det. Im. GD Alignment Inclusion
LIAR [47] 13K
DFIM-HQ [25] 140K
MEIR [36] 139k
MiRAGeNews [13] 15k
COSMOS [41] 453k
$\text{DGM}^{4}$ [38] 230k
@@
Figure 2: MDSM statistics. (a) Heat-map of the distribution of different media sources and manipulation categories. (b) Types of face attributes. (c) The distribution of the manipulated image area as a proportion of the entire image in the image modality.
$\triangleright$ Text Fabrication for Face Swap. We use Qwen2-VL [45] to generate consistent narratives matching the manipulated image $I_{fs}$. This requires knowing: 1) the inserted person’s name $N_{c}$, and 2) the names of people remaining in $I_{fs}$. We extract the full name list from the original text $T_{s}$ using BERT [8], then input these names and the face $F_{s}$ into Qwen2-VL to identify the name matching $F_{s}$. We replace this name with $N_{c}$ to form the final name set $N_{A}$. Finally, we input $I_{fs}$ and a prompt (as shown in Fig. 1c) into Qwen2-VL to generate aligned text.
$\triangleright$ Face Attribute Manipulation. Face emotion editing, is also considered in our dataset (Fig. 1b). Our pipeline uses StyleCLIP [30] and HFGI [46] for attribute manipulations. For image $I_{s}\in D_{s}$, we analyze facial expressions using an emotion classifier [26] to determine emotional attributes (positive or negative). We randomly select a method to manipulate the primary face $F_{s}$’s attributes inversely to the classification outcome, producing $I_{fa}$. To ensure diversity, we control manipulation intensity with varied emotion prompts and introduce age modifications. The distribution of face attribute prompts is shown in Fig. 2(b), with $y_{box}$ stored as annotation.
$\triangleright$ Text Fabrication for Face Attribute. Similar to face swapping, text forgery for face attribute editing is also generated by Qwen2-VL but with distinct prompts. Specifically, we instruct the MLLM to focus primarily on facial expressions while downplaying unchanged elements. The full name list is initially extracted from the source text $T_{s}$ and incorporated into the prompts (Fig. 1c) in a manner analogous to face swapping.

### v2

# Yuchen Zhang

@@ Abstract
The detection and grounding of multimedia manipulation has emerged as a critical challenge in combating AI-generated disinformation. While existing methods have made progress in recent years, we identify two fundamental limitations in current approaches: (1) Underestimation of MLLM-driven deception risk: prevailing techniques primarily address rule-based text manipulations, yet fail to account for sophisticated misinformation synthesized by multimodal large language models (MLLMs) that can dynamically generate semantically coherent, contextually plausible yet deceptive narratives conditioned on manipulated images; (2) Unrealistic misalignment artifacts: currently focused scenarios rely on artificially misaligned content that lacks semantic coherence, rendering them easily detectable. To address these gaps holistically, we propose a new adversarial pipeline that leverages MLLMs to generate high-risk disinformation. Our approach begins with constructing the MLLM-Driven Synthetic Multimodal (MDSM) dataset, where images are first altered using state-of-the-art editing techniques and then paired with MLLM-generated deceptive texts that maintain semantic consistency with the visual manipulations. Building upon this foundation, we present the Artifact-aware Manipulation Diagnosis via MLLM (AMD) framework featuring two key innovations: Artifact Pre-perception Encoding strategy and Manipulation-Oriented Reasoning, to tame MLLMs for the MDSM problem. Comprehensive experiments validate our framework’s superior generalization capabilities as a unified architecture for detecting MLLM-powered multimodal deceptions. In cross-domain testing on the MDSM dataset, AMD achieves the best average performance, with 88.18 ACC, 60.25 mAP, and 61.02 mIoU scores.
[…]
figure Comparison between widely-used $\text{DGM}^{4}$ [DGM4_TPAMI] and MDSM (ours). (left) $\text{DGM}^{4}$ typically treats visual context manipulation as two independent procedures as rule-based text editing and image editing. The absence of context integration often results in poorly aligned samples, which can be readily perceived by the public. (right) In a real-world scenario, we usually face well-wrapped content on purpose. To mimic such a situation, we propose a new manipulation paradigm, which explicitly provides the modified image as well as meta info of facial editing to Multimodal Large Language Model (MLLM). Then we harness MLLM to generate contextually consistent, deceptive texts to form the challenging image-text pairs.
@@ 1 Introduction
Generative techniques have become a transformative force in artificial intelligence [wu2024deepseekvl2, peng2025lmmr1, dawn2024_ICLR, fu2025speculative, Florence2_CVPR, zhang2025ctrl], showing remarkable adaptability across various domains and modalities. These advancements, while enriching multimedia content, also pose significant challenges to information security. In the media industry in particular, maliciously fake content manipulated by such models can profoundly mislead audiences [Defen_fake_news_NeurIPS, deepfake_eccv_shao]. The unchecked spread of fake media has already negatively affected political, financial, and other sectors [cantarella2023does_fakenews, fake_news_impact_financial, rocha2021impact_health], gradually becoming a major social issue [Olan2024_fakeNewsImpact].
@@
While various fake news scenarios have been explored, including [luo_newsclippings] investigation of out-of-context social image-text pairs and [DGM4] work on detecting randomly tampered regions or words. Our analysis reveals two critical limitations in existing research: 1) Neglect of emerging risks from MLLMs: Current paradigms predominantly address rule-based text manipulation, overlooking the sophisticated linguistic capabilities of modern multi-modal large language models (MLLMs). MLLM-generated text exhibits superior fluency and contextual coherence, significantly increasing deception potential and public susceptibility. 2) Semantic misaligned artifacts. Most methodologies independently manipulate visual and textual elements, producing semantically discordant multimedia outputs. This misalignment not only renders manufactured disinformation too simplistic to effectively deceive the public, but also fails to replicate real-world adversarial behavior, as sophisticated attackers typically maintain meticulous visual-textual consistency to maximize manipulative impact. Both limitations render the multi-modal disinformation scenarios considered in existing works insufficiently realistic.
@@
To address these weaknesses, we take MLLM into consideration and focus on detecting the semantic-aligned manipulation. We first construct the MLLM-Driven Synthetic Multi-modal (MDSM) dataset, which manipulates image and text in a coordinated fashion using MLLM. For the image manipulation, we consider the typical Face Swap and Face Attribute editing. For text, We innovatively guide MLLM to generate modality-aligned yet misleading fake narratives using image editing metadata. As shown in Fig. [Uncaptioned image] The Coherence Trap: When MLLM-Crafted Narratives Exploit Manipulated Visual Contexts(b), after replacing Donald Trump’s face with Micheál Martin’s, we use the swapped name, Micheál Martin, to guide MLLM in generating text, ensuring that the named entity in the text aligns with the image. Following this strategy, we construct over 441k sample pairs.
@@
The alignment of modalities and the authentic texts from MLLMs pose significant challenges for the detection of manipulated media. First, the strategy of perceiving inconsistencies between images and text through contrastive learning, as employed by prior works [zhang2024asap, DGM4], is ineffective in MDSM where images and text are well-matched already. Merely observing aligned image-text pairs is inadequate for reliable detection. Consequently, external clues and contextual knowledge are essential. Second, existing architectures like ASAP [zhang2024asap] and HAMMER [DGM4], which feature multiple detection and grounding heads, are complex and lack generalizability to unseen media sources. To address these challenges, we propose Artifact-aware Manipulation Diagnosis via MLLM (AMD), which leverages MLLMs’ comprehensive understanding of real-world multimedia and their ability to provide unified textual outputs. And AMD generates detection and grounding results in a coherent, text-based format, offering a more intuitive and generalized solution. In summary, we highlight our contributions of this paper as follows:
@@
  • Comprehensive evaluations validate AMD’s effectiveness and generalization capability, outperforming existing methods while maintaining parameter efficiency. With only 0.27B parameters, AMD achieves the best domain generalization average performance on both MDSM (88.18 ACC) and $\text{DGM}^{4}$ (74.47 ACC).
@@ 2 Related Work
Deepfake Detection. The rapid progress of generative models and the surge in synthetic content have accelerated advances in Deepfake detection. Existing work spans unimodal and multimodal approaches. Unimodal, including image-based [ICCV_9710718, CVPR_9157215] and text-based [ACL_NEP_text_unimodal, huang2023faking] approaches, already achieve strong results. With the rise of Multimodal Large Language Models (MLLMs), multimodal Deepfake detection has received increasing attention [liu2024mmfakebench, DGM4, liu2024fkaowl]. Regarding datasets, Shao et al. [DGM4] introduced the pioneering DGM4 benchmark for multimodal manipulation detection and grounding. However, its manipulations are rule-based, leading to semantically fragmented image-text discrepancies that do not accurately reflect real-world misinformation. MMFakeBench [liu2024mmfakebench] recognized this limitation and proposed generating semantically aligned news images using text-to-image models. Yet such semantically matched samples constitute only 30% of its fake subset, and the dataset contains merely 11k samples, limiting its utility for training robust detectors. Existing Deepfake datasets also fail to consider the risk of semantically coherent but misleading text generated by modern MLLMs. On the methodological side, HAMMER [DGM4] integrates contrastive learning to build a detector capable of classifying manipulation types and grounding manipulated regions, but it does not address cross-domain robustness. Beyond conventional multimodal detectors, FKA-Owl [liu2024fkaowl] employs a 7B-scale MLLM with several architectural modifications to enhance generalization. However, it is trained on DGM4, where text manipulations follow fixed editing rules rather than being synthesized by MLLMs, making it unsuitable for detecting more subtle, semantically aligned misinformation produced by modern models. Moreover, FKA-Owl performs only binary real/fake classification without fine-grained manipulation type prediction or grounding, and its large backbone and heavy architectural design result in substantially increased computational cost and slower inference.
@@
Multi-Modal Large Language Model. In recent years, Multi-Modal Large Language Models have emerged as a crucial technology for understanding and reasoning across multiple modalities, particularly text and images. By extending the capabilities of Large Language Models (LLMs) to incorporate visual inputs, these models have demonstrated outstanding performance in tasks such as image captioning and visual question answering. CLIP  [CLIP_Radford2021LearningTV] and ALIGN  [ALIGN_Jia2021ScalingUV] leveraged contrastive learning to align visual and textual representations, enabling efficient zero-shot vision-language understanding. Subsequently, models such as Flamingo  [Flamingo_NEURIPS2022] and BLIP-2  [BLIP2_ICML] have introduced vision-language transformers, integrating pre-trained LLMs with vision encoders to enhance cross-modal reasoning and generative capabilities. More recently, GPT-4V  [GPT4TR] and Florence-2  [Florence2_CVPR] have significantly enhanced the potential of MLLMs in tackling complex multi-modal tasks by leveraging a more efficient framework and larger-scale pre-training data. A key advantage of MLLMs is their acquisition of extensive world knowledge through large-scale pretraining, which substantially strengthens their reasoning abilities in downstream tasks. Such knowledge not only enhances cross-modal understanding but also proves essential for handling challenging problems, including misinformation detection.
@@ 3 MDSM Dataset Construction
Figure 1: MDSM Construction Pipeline. $\square$: The emotion detector helps the face editing model reverse the facial expression. The manipulated image is paired with MLLM-generated fabricated text for the Face Attribute&Text Fabrication category (FA&TF), or with authentic text for the Face Attribute category (FA) . $\square$: The MLLM fabricated text paired with authentic image for the Text Fabrication category (TF). $\square$: After swapping the face, the name list is updated for text-image alignment, and the manipulated image is paired with MLLM-generated fabricated text for the Face Swap&Text Fabrication category (FS&TF), or with aligned authentic text for the Face Sap category (FS).
As shown in Fig. 1, the collected source news data undergoes two key synthesis processes: 1) advanced image editing models generate visual manipulations, and 2) MLLMs produce text narratives that are semantically aligned with these visuals. We elaborate on these processes below.
@@ 3.1 Multi-Modal Media Source Collection
We use the GoodNews  [GoodNews], VisualNews  [VisualNews], and N24News  [N24News] datasets as the Source News Pool $\mathcal{D}_{O}$, which consists of over 2.1M image-text pairs sourced from various real-world news outlets. Given the significant influence of human-centric news among various forms of multi-modal media, we focus on human-centric data for MDSM. $\mathcal{D}_{O}$ is firstly filtered by detecting faces in images with Dlib  [dlib] and identifying person names in texts with BERT  [devlin2018bert]. Only pairs, $p_{s}=(I_{s},T_{s})$, containing both faces and named entities are used for manipulation. Additionally, we collect the Celebrity Head-shot Dataset $\mathcal{D}_{C}$, which contains about 30k pairs of head-shot images and corresponding names to facilitate the aligned manipulation for Face Swap. Details are provided in the appendix.
@@ 3.2 Multi-Modal Media Manipulation
In the image modality, two main attacks, Face Swap (FS) and Face Attribute (FA), are employed. For the text modality, we utilize advanced MLLM to generate semantic-aligned texts for the images.
Table 1: Comparison of the proposed MDSM with existing misinformation datasets, where $M$ Det., Text Det., Man. Type Det., and Im. GD stand for Multi-media Detection, Text Detection, Manipulation Type Detection and Image Grounding.
Datasets Samples Modality Tasks Training Semantic MLLM
Text Image MM Det. Text Det. Man. Type Det. Im. GD Support Alignment Inclusion
LIAR  [Text_news_dataset_liar] 13K
DFIM-HQ  [DFIM_HQ] 140K
MEIR  [MEIR_ACM18] 139k
MiRAGeNews  [huang-etal-2024-miragenews] 15k
COSMOS  [COSMOS_AAAI] 453k
MMFakeBench  [liu2024mmfakebench] 11k
$\text{DGM}^{4}$  [DGM4] 230k
@@
Figure 2: MDSM Statistics. (a) Distribution of media sources and manipulation categories. (b) Types of face attributes. (c) Distribution of the manipulated image area proportion of the entire image.
$\triangleright$ Face Swap. Face swap is a critical tool for attackers to forge images of public figures and politicians, posing threats to societal security. We use two representative face swap methods, SimSwap  [chen2020simswap] and e4s  [liu2022fine_e4s], to perform such manipulations. We prioritize modifying larger faces to target the primary subject in the image (Fig. 2c). Given a source image $I_{s}$, we randomly choose one of the two methods and replace the largest face $F_{s}$ in $I_{s}$ with a face $F_{c}$ from $\mathcal{D}_{C}$, generating a manipulated face swap sample $I_{fs}$. The bounding box $y_{box}=\{x_{1},y_{1},x_{2},y_{2}\}$ of the swapped face and the name of $F_{c}$ are recorded. To keep the image-text aligned, corresponding processing is also done on the authentic text. We use MLLM to identify the name of $F_{s}$, as shown in Fig. 1, $F_{s}$ is identified as Trump. We then refine the authentic text for FS category by replacing this name with $F_{c}$’s name.
$\triangleright$ Face Swap and Text Fabrication. We use Qwen2-VL  [Qwen2-VL] to generate consistent but misleading narratives. This requires knowing: 1) the inserted person’s name, and 2) the names of people remaining in $I_{fs}$. Using the same strategy as above, we get $F_{s}$’ name. We extract the full name list from the original text $T_{s}$ using BERT  [devlin2018bert], and then replace $F_{s}$ with inserted person’s name to form the final name list $N_{A}$. Finally, we input $I_{fs}$ and $N_{A}$ into Qwen2-VL to generate aligned text.
$\triangleright$ Face Attribute. Face emotion editing is also considered in our dataset. Our pipeline uses StyleCLIP  [patashnik2021styleclip] and HFGI  [wang2022high_HFGI] for attribute manipulations. Firstly, we analyze facial expressions using an emotion detector  [Arriaga2017RealtimeCN] to determine positive or negative emotions. We then randomly select a method to manipulate the primary face $F_{s}$’ attributes inversely to the classification outcome, producing $I_{fa}$. To ensure diversity, we control manipulation intensity with variable prompts and introduce age modifications. The distribution of face attribute prompts is shown in Fig. 2(b), with $y_{box}$ stored as annotation. Since the characters in $I_{fa}$ have not changed, the paired text in this category is still authentic.
$\triangleright$ Face Attribute and Text Fabrication. Similar to face swapping, text forgery for face attribute editing is also generated by Qwen2-VL but with distinct prompts. Specifically, we instruct the MLLM to focus primarily on facial expressions to generate narratives that conform to the characters’ demeanor. The input full name list is initially extracted from the source text $T_{s}$.
$\triangleright$ Text Fabrication. For the TF category, we also use BERT  [devlin2018bert] to extract the name list $N_{A}$ from the original text. Then, we input $N_{A}$ and the original image into the MLLM to generate narratives that match the implied meaning but are still fabricated.